# Design and Development of Portable RPC-Based Cosmic Muon Tracker

### v2

@@ 7 System Efficiency, Technical Challenges and Student Training
The Cosmic Muon Tracker (CMT) was evaluated extensively to assess detector performance, data acquisition stability, portability and its utility in student training. The primary metric of interest was the efficiency of the Resistive Plate Chambers (RPCs), determined from muon track reconstruction. Muon tracks were fitted using a least-square method to estimate expected hit positions across each RPC layer. The measured hits were then compared against the expected trajectory to compute efficiency. Most layers demonstrated high efficiency, with six out of eight RPCs exceeding (90%) as summarized in Table 1. Instances of reduced efficiency during field deployment prompted RPC repair or replacement to maintain system reliability.
@@
The LED event display integrated into the Data Acquisition (DAQ) logic provides real-time visual feedback by illuminating LEDs corresponding to muon hits Figure 18. Its stability under varying environmental conditions such as temperature, humidity and pressure was closely monitored. These parameters influence the RPC noise rate and signal multiplicity, thereby impacting the clarity of track visualization. Continuous logging of strip noise rates allowed the detection of abnormalities and ensured consistent detector performance during long-term operation.
The CMT system was also deployed as a platform for student engagement in hands-on detector development and outreach as shown in Figure 19. Graduate and undergraduate students were trained in operating the DAQ system, troubleshooting noise issues, operating RPC voltages, and analyzing muon data. This practical exposure to high-energy physics instrumentation significantly contributed to technical skill development.
CMT serves as a scalable and portable platform for RPC detector development and student-led particle physics research as well as used in INO-ICAL project outreach programs such as Vigyan Samagam.
@@ 8 Conclusion
CMT is a gaseous detector that requires regular monitoring to ensure optimal performance. The tracker is flushed with fresh gas mixture whenever its performance degrades. Its performance is also tested under various ambient conditions. The compact and portable design of the CMT makes it ideal for measuring muon flux at different altitudes. With a timing accuracy of , it can determine the direction of muon paths and thus can be used for tomography. It can serve as a simple counter for background measurements in mines, etc. The CMT can also function in both vertical and horizontal orientations allowing for comparisons of horizontal and vertical fluxes at different locations as well as any zenith angle.

### v3

@@ 6 System Efficiency, Technical Challenges and Student Training
CMT was evaluated extensively to assess detector performance, data acquisition stability, portability and its utility in student training. The primary metric of interest was the efficiency of the RPCs, determined from muon track reconstruction. Muon tracks were fitted using a least-square method to estimate expected hit positions across each RPC layer. The measured hits were then compared against the expected trajectory to compute efficiency. Table 1 summarizes the measured efficiency for each RPC layer.
@@
While two layers exceeded 96% efficiency, several others showed lower values, ranging between 50–90%. Notably, some of the lower-efficiency chambers had shown better performance during initial cosmic-ray testing prior to stack assembly. This suggests that performance degradation may have occurred during handling or post-integration. These deviations are due to multiple contributing factors including minor gas leakage leading to degraded avalanche formation and post-assembly mechanical stress. In addition, the large differences in efficiency arose primarily from layer-specific hardware conditions such as electrode contact issues, variation in electronic threshold settings and local noise pickup in certain readout channels. For long-term operation of the muon telescope, degradation in efficiency and stability is unacceptable. These issues are presently mitigated through periodic maintenance, including inspection and repair of interconnects, along with regular flushing of a fresh gas mixture at least once every five days or whenever a decline in performance is detected. To support this procedure, a pre-mixed 20-liter gas cylinder is permanently assigned to the system and transported with the stack during deployments. Also to ensure all RPCs operate in high and stable efficiency regions for precise muon flux measurements, future revisions of the CMT stack will include improved gas sealing, mechanical reinforcement to minimize stress and contact inconsistencies, and efficiency stability tests conducted before and after each deployment. We will also use high-quality interconnects for handling RPC signals and high voltages, along with proper grounding of the front-end electronics and DAQ systems. Additionally, adopting a more modular stack design will enable in-situ testing of individual RPCs, provide increased space for debugging, and facilitate easier removal and replacement of RPC units.
The LED event display integrated into the DAQ logic provides real-time visual feedback by illuminating LEDs corresponding to muon hits [15], as shown in Figure 18. Its stability under varying environmental conditions such as temperature, humidity and pressure was closely monitored. These parameters influence the RPC noise rate and signal multiplicity, thereby impacting the clarity of track visualization. The CMT system was also deployed as a platform for student engagement in hands-on detector development and outreach programs such as Vigyan Samagam as shown in Figure 19.
@@ 7 Conclusion
CMT is a portable, gas-based detector developed primarily for outreach, educational demonstrations and simple laboratory or field measurements of cosmic ray muons. The compact and portable design of the CMT makes it ideal for measuring muon flux at different altitudes. With a timing accuracy of , it can determine the direction of muon paths and thus can be used for tomography. It can serve as a simple counter for background measurements in mines, etc. The CMT can also function in both vertical and horizontal orientations allowing for comparisons of horizontal and vertical fluxes at different locations as well as any zenith angle. Overall, the CMT offers a reliable and accessible platform for introducing students and the public to particle detection concepts, while also supporting preliminary muon flux measurements and background radiation studies.